OpenStudy (anonymous):

What is the solution to the system of equations? 2x – y = 7 y = 2x + 3

OpenStudy (anonymous):

lol yup :)

OpenStudy (anonymous):

Hey i forgot how to solve this so can u plz help :)

pooja195 (pooja195):

use substituion

pooja195 (pooja195):

\[2x – (2x+3) = 7\]

pooja195 (pooja195):

\[–1(2x+3) \] distribute

OpenStudy (anonymous):

oh yeah now i remember thanks a lot

pooja195 (pooja195):

lol your welcome would u like to work it out? Or do u get it

OpenStudy (anonymous):

No can we work it out plz ?

pooja195 (pooja195):

:)

pooja195 (pooja195):

This is what we get when we substitute y in to equation 1 2x–(2x+3)=7 We need to distribute –1(2x+3) what do you get once you distribute?

OpenStudy (anonymous):

-2x+-3

pooja195 (pooja195):

good so we have.... \[2x–2x+3=7\] Combine the like terms

OpenStudy (anonymous):

3=7

pooja195 (pooja195):

no

pooja195 (pooja195):

well yes but 0x+3=7 (we need to keep 0)

OpenStudy (anonymous):



pooja195 (pooja195):

wait there is no solution :/

OpenStudy (anonymous):

oh ok that makes sense now THANKS a lot i remember this stuff now :)

pooja195 (pooja195):

haha no probs :)